OpenStudy (anonymous):

replace the ___ with <,>, or = to make a true statement 7^4/7^2 ________ 49

OpenStudy (anonymous):

Hi! Need help?

OpenStudy (anonymous):

yes please @AMYCARTER

OpenStudy (kendricklamar2014):

1. We need to solve: 7^4/7^2

OpenStudy (anonymous):

< is the answer because 28/14 is 2<49

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

yes @cherry18

OpenStudy (kendricklamar2014):

PLease @cherry18 and @PizzaLover123 show ur work so @princess_51 know how to solve this equation

OpenStudy (anonymous):

Cherry did @KendrickLamar2014

OpenStudy (anonymous):

I did @KendrickLamar2014 by saying 28/14

OpenStudy (anonymous):

So there

OpenStudy (kendricklamar2014):

All of it?

OpenStudy (anonymous):

Great job Cherry! :) @cherry18 Ignore him xD

OpenStudy (kendricklamar2014):

Did you tell @princess_51 how you got 28/14?

OpenStudy (anonymous):

yeah

OpenStudy (kendricklamar2014):

and by the way < is not the answer

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}7^4 \\ 7^2\end{matrix}\right) = 49\]

OpenStudy (kendricklamar2014):

\[49 = 49\]\ Answer: = @princess_51

OpenStudy (anonymous):

NO ITS NOT

OpenStudy (kendricklamar2014):

Yes it is

OpenStudy (anonymous):

Nope

OpenStudy (kendricklamar2014):

Yup \[7^4/7^2 = 49\]

OpenStudy (kendricklamar2014):

\[7^4 is~the~same~as~ 7*7*7*7=2401\] \[7^2~is~the~same~as~7*7 = 49\] \[\left(\begin{matrix}2401 \\ 49\end{matrix}\right) = 49\]

OpenStudy (anonymous):

I got sorry.

OpenStudy (kendricklamar2014):

It's ok everybody makes mistakes sometimes :)